# Utilizing anticoincidence veto in a search for gravitational-wave transients

### v1

@@ Abstract
We devise a technique to suppress the effect of noise transients occurring at gravitational-wave detectors based on temporal anticoincidence. Searches for gravitational-wave signals in the detector data are prone to spurious disturbances of terrestrial origin. The technique presented here benefits from the fact that such effects are generally non-coincident in time at geographically separated detectors. Therefore, abnormally loud triggers that are not time-coincident can be vetoed. We implement the veto technique in a matched-filter search for transient signals from binary black holes and observe search backgrounds to be generally close to the Gaussian limit. An improvement in the sensitivity of the search is demonstrated using simulated signals. The technique is expected to especially improve the detection efficiency of the search toward short duration transient signals.
@@ I Introduction
In the observing mode, the output of a detector primarily comprises a stationary Gaussian noise floor. It also contains instances of abrupt disturbances of instrumental or environmental origin, local to the given detector. In addition to these terrestrial disturbances, the detector output contains occasional instances of transient astrophysical GW signals. Such signals are expected to be coincident in time across all observing detectors. Therefore, coincident observations enhance the chances for confident detection of a GW candidate signal. Searches for the candidate GW signals from the detector data involve multiple steps. The presence of noise transients degrades the sensitivity of the searches toward the GW transients. The sensitivity of a search can further depend upon several other factors, including the response of the search to the noise transients.
@@
Primary aligned-spin, $\mathrm{\chi_{2z}}$ [-0.99, 0.99]
@@
The general techniques used in the search are outlined as follows. The search relies on matched-filtering the data from the detectors (Sathyaprakash and Dhurandhar, 1991). It uses a template waveform that models the incoming GW signal, which is generated using the IMRPhenomXAS model, computed directly in the frequency domain (Pratten et al., 2020). The power spectral density (PSD) computed from the detector data is used which is obtained following Welch’s method (Welch, 1967). A matched-filter results into a signal-to-noise ratio (SNR) timeseries per detector from which triggers are generated by thresholding (Usman et al., 2016). The SNR is reweighted using the power $\mathrm{\chi\textsuperscript{2}}$-veto (Allen, 2005), followed by the sine-Gaussian veto applicable for short duration templates (Nitz, 2018). We select a duration of 30 days during the first half of the third observing run (O3a) of the Advanced LIGO (Abbott et al., 2023b). Simulated signals are injected into the data and the PSO search is used to recover the signals (Pal and Nayak, 2024). To estimate a statistical significance to a candidate event, the time-sliding approach is used (Wąs et al., 2009). The search background and other aspects of the search are investigated with and without using the anticoincidence veto introduced in this work.
The article is organized into the following sections. Section II describes the process used to set up the search. We generally skip the detailed description of the computational steps as they are reported elsewhere as mentioned earlier. Section III describes the implementation of the anticoincidence veto in the search. Comparisons of various aspects of the search with and without the veto are included. Section IV describes an assessment for the detection efficiency of the search using the injected signals. A few aspects of the search background involving higher-mass systems and higher rate of signal detection are discussed in the appendix sections.
@@ II Search description
Here we continue by assuming that the algorithm furnishes an optimized template point as the output from any given coincident data stretch. The secondary triggers, which are computed from the optimized template in the same fashion as described above, undergo the same veto processes as the primary triggers. The collection of these triggers over all data stretches is depicted in Fig. 2. These are additionally tested for time-coincidence (vetoed and foreground events generated) and are used to assign a statistical significance to the generated foreground events, as described below. Here a 15 ms window is used for the HL detector network, given the travel time of light between the two detectors. Sufficiently loud triggers of astrophysical origin are expected to pass the time-coincidence test and feature as foreground events, whereas any non-Gaussian triggers of terrestrial origin are likely to pass a time-anticoincidence test as described in the next section. Note that the only point of contact between the data and the search is through the optimized template obtained from the data stretches.
@@
The triggers obtained during the buffer time from the detectors are artificially time-slided by an interval greater than the light travel time between the detectors. Such unphysical time-slide duration ($\sim$ 50 ms) ensures that any coincidence between triggers hence formed are non-astrophysical in nature. The false alarm rate (FAR, or equivalently the significance) is assigned to a foreground candidate based on the number of such background events having a ranking statistic larger than the given foreground event. A large sample of background events is generated with as many time-slides as possible, which is limited by the number of distinct time-slides possible with a given buffer duration. When the procedure described above assigns acceptable FAR values to foreground events and an appropriate number of foreground events are generated using the triggering process described earlier, we observe an empirical self-consistency, as shown in Fig. 4.
In the O3-like scenario, the entire analysis requires modest computing resources ($\sim$ 8 CPU cores and $\sim$ 8 GB memory) to finish the computation for a given stretch of coincident data in time before the next stretch is analyzed. Triggers from only the optimized templates are stored on the computer disk, which occupy $\sim$ 10 GB for a full typical observing run.
@@ III Anticoincidence veto
An astrophysical trigger in any detector should be accompanied with a coincident counterpart during a coincident observing time [^1] There are exceptions to this, e.g., when the source is located around a blind spot arising due to antenna pattern functions of a detector, or due to a reduced sensitivity of the detector during observations; such cases are expected to be small in number.. Therefore, the triggers in excess can be effectively vetoed. We devise a simple technique to veto such triggers based on anticoincidence. Triggers in a given detector that are louder than a certain threshold (of $\sim$ 6 in the reweighted SNR) are tested for time-anticoincidence with any nearby trigger from the remaining detector. If the test is positive, the veto is turned on and the triggers are removed, effectively masking the trigger streams for a short duration as described below. When the test is negative, the triggers from each detector are retained. A summary of this logic is shown in Table 2. We observe that the distribution of the resulting triggers after the veto process is close to the Gaussian case in both the detectors. Note that there are a few loud surviving triggers in the detectors which fail the anticoincidence test, i.e., they pass the coincidence test and can be potentially astrophysical in origin. Broadly, the veto technique thus reduces the non-Gaussian and non-stationary component of the instrument data.
@@
Since the time-coincidence test requires a trivial computation, the overall cost for the application of the veto is negligibly small. For instance, this one-time step can be completed within a few seconds on a single CPU for a typical observing run.
@@ IV Detection efficiency
In this section, we demonstrate the impact of the veto technique on the sensitivity of the search. To evaluate the sensitivity, simulated signals are injected into real instrument data and are recovered by the search with and without the veto applied. The signal parameters come from the injection sets containing a representative population of the astrophysical sources (o3p, ). The injections are pre-selected from the coincident duration used earlier, and are bounded by the signal parameter ranges given in Table 1. In addition to the aligned-spin components, the injections also carry precessing spins, implemented using the IMRPhenomXP waveform model (Pratten et al., 2021).
A search is conducted on the triggers, including those due to the injections, using the same analysis configuration as described in the earlier sections. Similar to the analysis framework, the triggers from both detectors are channeled through two different processes- with and without the veto step applied. Foreground candidates are computed from the triggers and a FAR estimate is assigned to each candidate. Triggers around the injection times are properly accounted for in order to prevent them from affecting the FAR estimates of any foreground candidate. The injections are identified from the foreground events as follows. An injection is considered to be recovered if there exists a foreground candidate with triggers at the two detectors appearing in coincidence within 1 second around the injected coalescence time above a significance threshold. The number of the recovered injections up to a given significance threshold (in terms of inverse FAR or IFAR) with and without the veto step applied is summarized in Fig. 6. We observe that the number injections recovered with the veto applied is greater than that without the veto, at any of the given IFAR thresholds. This is attributed to the improved backgrounds with the veto as indicated in Fig. 5.
@@ V Conclusions
Here we have used only two detectors for demonstration and have not considered any possible single detector events in the coincident times. A search including single detector events can be explored in the future which could utilize the byproducts of this search. On the other hand, the technique can also be extended to a larger detector network especially with comparable sensitivities.
@@ VI Acknowledgments
The research and development was carried out at the Center of Excellence in Space Sciences India (CESSI). CESSI is a multi-institutional Center of Excellence hosted by the Indian Institute of Science Education and Research (IISER) Kolkata and has been established through funding from the Ministry of Education, Government of India. We acknowledge the use of IUCAA LDG cluster Sarathi for the computational/numerical work. Several tools from the PyCBC software package are used in the analyses. We acknowledge interaction and feedback from several members of the GW community. The author thanks K. Rajesh Nayak for the discussions during this work. The author thanks Johann Fernandes for carefully reading the manuscript. This material has a LIGO document number, P2500796. The author did not receive funding or financial support from any agency or institution.
@@ VII Data availability
Appendix A Searching for signals from higher-mass systems
Dedicated searches for intermediate mass black hole binaries have been successfully carried out (Abbott et al., 2019b, 2022; Chandra et al., 2021). In this section, we implement a search similar to the one indicated in Table 1, except that the targeted component masses now belong to $[100,250]$ $\mathrm{M_{\odot}}$. This further exposes the search to the possible glitches triggered by relatively higher-mass (shorter duration) templates. The sensitivity of a search can be optimized by using a strict cut on the duration of templates, see for example (Chandra et al., 2021). Here we skip such fine optimization choices, primarily to isolate the effectiveness of the veto step.
The injections within the given mass range are recovered from a longer duration of coincident data during O3a in a similar fashion as described in the main text. These are summarized in Fig. 8. We observe that the number of injections recovered is greater while using the anticoincidence veto at any given significance threshold. The self-consistency of the significance estimates is also shown. We attribute the greater improvement in the injection recovery, in part, to the smaller effect of the joint injections within any analysis segment as described in Section IV, which is around 4% here due to a smaller rate of the high-mass injections in the given dataset, with the same analysis configuration. We note that the computational cost of conducting such a search is trivially small.
@@ Appendix B Significance estimates in the presence of nearby coincident events
In the earlier sections, we discussed the effect of anticoincident triggers on searches for transient signals. However, nearby (loud) coincident foreground triggers of noise origin or of true astrophysical origin, to a given candidate can still limit its significance estimate. This is illustrated for GW190521_030229 (referred to as GW190521) in Fig. 9. The kink in the inclusive significance estimate is due to a coincident candidate at the GPS time 1241858977.0 seconds, around 80 hours backward in coincident time from the event, obtained from the same search. This is a marginal candidate with an IFAR of $\sim$ 1.0 year and with a network SNR of 10.65, which inhibits the significance estimate beyond a certain value of $\sim$ 50000 years. The coincident events obtained from a few blocks of data around GW190521 are summarized in Table 3. We do not observe such an effect from the confident event GW190519_153544 (referred to as GW190519), the nearest neighbour to GW190521 backward in time, and other remaining candidates, due to their relatively smaller detector-wise ranking statistics. Though very high significance is typically not necessary for the observation of astrophysical candidates, such cases are expected to be critical in the upcoming observations given a higher projected rate of detections. To take this into account, an appropriate workaround to the time-slide approach may be needed in the future.
@@
Table 3: Summary of the candidate events from the search. Here events with an IFAR less than 100 years and more than or equal to 1 year are typed marginal. Reported IFARs are all-inclusive. Not all marginal events may be reported elsewhere.

### v2

@@ Abstract
We devise a technique to suppress the effect of noise transients occurring at gravitational-wave detectors based on temporal anticoincidence. Searches for gravitational-wave signals in the detector data are prone to spurious disturbances of terrestrial origin. The technique presented here benefits from the fact that the noise effects are generally non-coincident in time at geographically separated detectors. Therefore, abnormally loud detector triggers that are not time-coincident can be vetoed. We implement the veto technique in a matched-filter search for transient signals from binary black holes and observe search backgrounds to be generally close to the Gaussian limit. An improvement in the sensitivity of the search is demonstrated using simulated signals. The technique is expected to especially improve the detection efficiency of the search for short duration gravitational waves.
@@ I Introduction
Searches for the candidate GW signals from the detector data involve multiple steps. In the observing mode, the output of a detector primarily comprises a stationary Gaussian noise floor. It also contains instances of abrupt disturbances of instrumental or environmental origin, local to a given detector. In addition to these terrestrial disturbances, the detector output contains occasional instances of transient astrophysical GW signals. Such signals are expected to be coincident in time across all observing detectors. Therefore, coincident observations generally enhance the chances for confident detection of a GW candidate signal. The presence of noise transients however, degrades the sensitivity of the searches toward the GW transients. The sensitivity of a search can depend upon several factors, including the response of the search to the noise transients.
@@
Primary aligned-spin, $\mathrm{\chi_{1z}}$ [-0.99, 0.99]
@@
The general techniques used in the search are outlined as follows. The search relies on matched-filtering the data from the detectors (Sathyaprakash and Dhurandhar, 1991). It uses a template waveform that models the incoming GW signal, which is generated using the IMRPhenomXAS model, computed directly in the frequency domain (Pratten et al., 2020). The power spectral density (PSD) computed from the detector data is used which is obtained following Welch’s method (Welch, 1967). A matched-filter results into a signal-to-noise ratio (SNR) timeseries per detector from which triggers are generated by thresholding (Usman et al., 2016). The SNR is reweighted using the power $\mathrm{\chi\textsuperscript{2}}$-veto (Allen, 2005), followed by the sine-Gaussian veto applicable for short duration templates (Nitz, 2018). We select a duration of 30 days during the first half of the third observing run (O3a) of the Advanced LIGO (Abbott et al., 2023b). Simulated signals are injected into the data and the PSO search is used to recover the signals (Pal and Nayak, 2024). To assign a statistical significance to a candidate event, the time-sliding approach is used (Wąs et al., 2009). The search background and other aspects of the search are investigated with and without using the anticoincidence veto introduced in this work.
The article is organized into the following sections. Section II describes the process used to set up the search. We generally skip the detailed description of the computational steps since they are reported elsewhere as mentioned earlier. Section III describes the implementation of the anticoincidence veto in the search. Comparisons of several aspects of the search with and without the veto are included. Section IV describes an assessment of the detection efficiency of the search using the injected signals. A few aspects of the search background involving higher-mass systems and higher rate of signal detection are discussed in the appendix sections.
@@ II Search description
Here we continue by assuming that the algorithm furnishes an optimized template point as the output from any given coincident data stretch. The secondary triggers, which are computed from the optimized template in the same fashion as described above, undergo the same veto processes as the primary triggers. The collection of these triggers over all data stretches is depicted in Fig. 2. These are additionally tested for time-coincidence (vetoed and foreground events generated) and are used to assign a statistical significance to the generated foreground events, as described below. Here a 15 ms window is used for the HL detector network, given the travel time of light between the two detectors. Sufficiently loud triggers of astrophysical origin are expected to pass the time-coincidence test and feature as foreground events, whereas any non-Gaussian triggers of terrestrial origin are likely to pass a time-anticoincidence test as described in the next section. Note that the only point of contact between the data and the search is through the optimized templates obtained from the data stretches.
@@
The triggers obtained during the buffer time from the detectors are artificially time-slided by an interval greater than the light travel time between the detectors. An unphysical time-slide duration (of $\sim$ 50 ms) ensures that any coincidence between triggers hence formed are non-astrophysical in nature. The false alarm rate (FAR, or equivalently the significance or the inverse FAR) is assigned to a foreground candidate based on the number of such background events having a ranking statistic larger than the given foreground event. A large sample of background events is generated with as many time-slides as possible, which is limited by the number of distinct time-slides possible with a given buffer duration. When the procedure described above assigns acceptable FAR values to foreground events and an appropriate number of foreground events are generated using the triggering process described earlier, we observe an empirical self-consistency, as shown in Fig. 4.
In O3-like scenarios, the entire analysis requires modest computing resources ($\sim$ 8 CPU cores and $\sim$ 8 GB memory) to finish the computation for a given stretch of coincident data in time before the next stretch is analyzed. Triggers from only the optimized templates are stored on the computer disk, which occupy $\sim$ 10 GB for a full typical observing run.
@@ III Anticoincidence veto
An astrophysical trigger in any detector should be accompanied with a coincident counterpart during a coincident observing time [^1] There are exceptions to this, e.g., when the source is located around a blind spot of a detector arising due to antenna pattern functions, or due to a reduced sensitivity of the detector during observations; such cases are expected to be small in number.. Therefore, the triggers in excess can be effectively vetoed. We devise a simple technique to veto such triggers based on anticoincidence. Triggers in a given detector that are louder than a certain threshold (of $\sim$ 6 in the reweighted SNR) are tested for time-anticoincidence with any nearby trigger from the remaining detector(s). If the test is positive, the veto is turned on and the triggers are removed, effectively masking the trigger streams for a short duration as described below. When the test is negative, the triggers from each detector are retained. A summary of this logic is shown in Table 2. We observe that the distribution of the resulting triggers after the veto process is close to the Gaussian case in both the detectors. Note that there are a few loud surviving triggers in the detectors which fail the anticoincidence test, i.e., they pass the coincidence test and can be potentially astrophysical in origin. Broadly, the veto technique thus reduces the non-Gaussian and non-stationary component of the instrument data.
@@
Since the time-coincidence test requires a trivial computation, the overall cost for the application of the veto is negligibly small. For instance, this one-time step can be completed within a few seconds on a single CPU for the entirety of a typical observing run.
@@ IV Detection efficiency
In this section, we demonstrate the impact of the veto technique on the sensitivity of the search. To evaluate the sensitivity, simulated signals are injected into real instrument data and are recovered by the search with and without the veto applied. The signal parameters come from the injection sets containing a representative population of the astrophysical sources (o3p, ). The injections are pre-selected from the coincident duration used earlier, and are bounded by the signal parameter ranges given in Table 1. In addition to the aligned-spin components, the injections also carry precessing spins, implemented using the IMRPhenomXP waveform approximant (Pratten et al., 2021).
A search is conducted on the triggers, including those due to the injections, using the same analysis configuration as described in the earlier sections. Similar to the analysis framework, the triggers from both detectors are channeled through two different processes- with and without the veto step applied. For each process, foreground candidates are computed from the triggers and a FAR estimate is assigned to each candidate. Triggers around the injection times are properly accounted for in order to prevent them from affecting the FAR estimates of any foreground candidate. The injections are identified from the foreground events as follows. An injection is considered to be recovered if there exists a foreground candidate with triggers at the two detectors appearing in coincidence within 1 second around the injected coalescence time above a significance threshold. The number of the recovered injections up to a given significance threshold (IFAR) with and without the veto step applied is summarized in Fig. 6. We observe that the number injections recovered with the veto applied is greater than that without the veto, at any of the given IFAR thresholds. This is attributed to the improved backgrounds with the veto as indicated in Fig. 5.
@@ V Conclusions
Here we have used only two detectors for demonstration and have not considered any possible single-detector events in the coincident times. A search including single detector events can be explored in the future which could utilize the byproduct triggers of this search. On the other hand, the technique can also be extended to a larger network of detectors especially with comparable sensitivities.
@@ VI Acknowledgments
The research and development was carried out at the Center of Excellence in Space Sciences India (CESSI). CESSI is a multi-institutional Center of Excellence hosted by the Indian Institute of Science Education and Research (IISER) Kolkata and has been established through funding from the Ministry of Education, Government of India. We acknowledge the use of IUCAA LDG cluster Sarathi for the computational/numerical work. Several tools from the PyCBC software package are used in the analyses. We acknowledge interaction and feedback from several members of the GW community. The author thanks K. Rajesh Nayak for the discussions during this work. The author thanks Johann Fernandes for carefully reading the manuscript and for providing feedback on it. This material has a LIGO document number, P2500796. The author did not receive funding or financial support from any agency or institution.
@@ VII Data availability
Appendix A Searches for higher-mass systems
Dedicated searches for intermediate mass black hole binaries have been previously carried out (Abbott et al., 2019b, 2022; Chandra et al., 2021). In this section, we implement a search similar to the one indicated in Table 1, except that the targeted component masses now belong to $[100,250]$ $\mathrm{M_{\odot}}$. This further exposes the search to the possible glitches triggered by relatively higher-mass (shorter duration) templates. The sensitivity of a search can be optimized by using, for example, a strict cut on the duration of templates (Chandra et al., 2021). Here we skip such fine optimization choices, primarily to isolate the effectiveness of the veto step.
The injections within the given mass range are recovered from a longer duration of coincident data during O3a in a similar fashion as described in the main text. These are summarized in Fig. 8. We observe that the number of injections recovered is greater while using the anticoincidence veto at any given significance threshold. The self-consistency of the significance estimates is also shown. We attribute the greater improvement in the significance estimates, in part, to the smaller effect of the joint injections within any analysis segment as described in Section IV, which is around 4% here due to a smaller rate of the high-mass injections in the given dataset, with the same analysis configuration. We note that the computational cost of conducting such a search is trivially small.
@@ Appendix B Significance estimates in the presence of nearby coincident events
In the earlier sections, we discussed the effect of anticoincident triggers on searches for transient signals. However, nearby (loud) coincident foreground triggers of noise origin or of true astrophysical origin, to a given candidate can still limit its significance estimate. This is illustrated for GW190521_030229 (referred to as GW190521) in Fig. 9. The kink in the inclusive significance estimate is due to a coincident candidate at the GPS time 1241858977.0 seconds, around 80 hours backward in coincident time from the event, obtained from the same search. This is a marginal candidate with an IFAR of $\sim$ 1.0 year and with a network SNR of 10.65, which inhibits the significance estimate beyond a certain value of $\sim$ 50000 years. The coincident events obtained from a few blocks of data around GW190521 are summarized in Table 3. Here we do not observe such an effect from the confident event GW190519_153544 (referred to as GW190519), the nearest neighbour to GW190521 backward in time, and other remaining candidates, due to their relatively smaller detector-wise ranking statistics. Although very high significance is typically not necessary for the observation of astrophysical candidates, cases such as the one described above are expected to be critical in the upcoming observations given a higher projected rate of detections. To take this into account, an appropriate workaround to the time-slide approach may be needed in the future.
@@
Table 3: Summary of the candidate events from the search. Reported IFARs are all-inclusive. Here events with an IFAR less than 100 years and more than or equal to 1 year are typed marginal. Not all marginal events may be reported elsewhere.